Compute the following using the residue theorem (complex analysis):

\int_{0}^{2\pi} \frac{1}{1+cos^{2}\theta} d\theta
Answer #1

2TT I do 1 + cos²o using residue Theorem lo e Coso = + e 2 let Z=elo z tz 2 dz= i elo do 10,0) unit circle with control dz= i- Od 2 du ilu²+6u+1) becouse 2 times unit circle. De u du 1 (u²+ butl) Now poles u2+ 64+1 u= -6 + 536-4 2 u= - 6 t 32 2 u = -